# Score identity Distillation: Exponentially Fast Distillation of Pretrained Diffusion Models for One-Step Generation

### v1

@@ Appendix B Analytic study of the toy example
$f_{\psi}(\bm{x}_{t},t)=\bm{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\psi\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}},$
which means $\psi^{*}=\theta$, then by Equation˜12 we have
${\delta}_{\phi,\psi}(\bm{x}_{t})=\sigma_{t}^{-2}(f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t))=-\frac{\psi}{1+\sigma_{t}^{2}}.$ (35)
@@
$\hat{L}_{\theta}^{(1)}={\delta}_{\phi,\psi}(\bm{x}_{t})^{T}{\delta}_{\phi,\psi}(\bm{x}_{t})=\frac{\psi^{2}}{(1+\sigma_{t}^{2})^{2}}$ (36)
Therefore, while $\hat{L}_{\theta}=\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})=\frac{\theta^{2}}{(1+\sigma_{t}^{2})^{2}}$ would provide useful gradient to learn $\theta$, its naive approximation $\hat{L}_{\theta}^{(1)}$ could fail to provide meaningful gradient.
@@
$\hat{L}_{\theta}^{(2)}$ $=\hat{L}_{\theta}^{(1)}+\frac{{\delta}_{\phi,\psi}(\bm{x}_{t})^{T}(f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g})}{\sigma_{t}^{2}}$
$=\frac{\psi^{2}}{(1+\sigma_{t}^{2})^{2}}-\frac{\psi}{\sigma_{t}^{2}(1+\sigma_{t}^{2})}(\bm{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\psi\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}-\bm{x}_{g})$
$=\frac{\psi}{(1+\sigma_{t}^{2})^{2}}\left[\bm{x}_{g}-\frac{\epsilon_{t}}{\sigma_{t}}\right].$
@@
$=-\frac{1}{1+\sigma_{t}^{2}}{\delta}_{\phi,\psi}(\bm{x}_{t})\nabla_{\theta}G_{\theta}(z)$
$\approx\frac{1}{1+\sigma_{t}^{2}}[\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})-S_{\phi}(\bm{x}_{t})]\nabla_{\theta}G_{\theta}(z).$
Appendix C Training and Evaluation Details and Additional Results.
@@
Learning rate of ($\psi$, $\theta$) 1e-5 5e-6 1e-5 5e-6
@@
Figure 9: Analogous plot to Fig. 6 for ImageNet 64x64, where the batch size for SiD is 1024 and learning rate is 5e-6. The FID declines fast until it suddenly diverges. Increasing the batch size to 8192 and lowering the learning rate to 4e-6, as shown in Fig. 6, has alleviated the issue of sudden divergence.

### v2

@@ Appendix D Analytic study of the toy example
$f_{\psi}(\boldsymbol{x}_{t},t)=\boldsymbol{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\psi\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}},$
which means $\psi^{*}=\theta$, then by Equation 12 we have
${\delta}_{\phi,\psi}(\boldsymbol{x}_{t})=\sigma_{t}^{-2}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t))=-\frac{\psi}{1+\sigma_{t}^{2}}.$ (35)
@@
$\hat{L}_{\theta}^{(1)}={\delta}_{\phi,\psi}(\boldsymbol{x}_{t})^{T}{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})=\frac{\psi^{2}}{(1+\sigma_{t}^{2})^{2}}$ (36)
Therefore, while $\hat{L}_{\theta}=\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})=\frac{\theta^{2}}{(1+\sigma_{t}^{2})^{2}}$ would provide useful gradient to learn $\theta$, its naive approximation $\hat{L}_{\theta}^{(1)}$ could fail to provide meaningful gradient.
@@
$\hat{L}_{\theta}^{(2)}$ $=\hat{L}_{\theta}^{(1)}+\frac{{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}{\sigma_{t}^{2}}$
$=\frac{\psi^{2}}{(1+\sigma_{t}^{2})^{2}}-\frac{\psi}{\sigma_{t}^{2}(1+\sigma_{t}^{2})}(\boldsymbol{x}_{t}\frac{1}{1+\sigma_{t}^{2}}+\psi\frac{\sigma_{t}^{2}}{1+\sigma_{t}^{2}}-\boldsymbol{x}_{g})$
$=\frac{\psi}{(1+\sigma_{t}^{2})^{2}}\left[\boldsymbol{x}_{g}-\frac{\epsilon_{t}}{\sigma_{t}}\right].$
@@
$=-\frac{1}{1+\sigma_{t}^{2}}{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})\nabla_{\theta}G_{\theta}(z)$
$\approx\frac{1}{1+\sigma_{t}^{2}}[\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})-S_{\phi}(\boldsymbol{x}_{t})]\nabla_{\theta}G_{\theta}(z).$
Appendix E Training and Evaluation Details and Additional Results.
@@
Learning rate of ($\psi$, $\theta$) 1e-5 4e-6 1e-5 5e-6
@@
[…]
figureAnalogous plot to Fig. 5 for ImageNet 64x64, where the batch size for SiD is 1024 and learning rate is 5e-6. The FID declines fast until it suddenly diverges. Increasing the batch size to 8192 and lowering the learning rate to 4e-6, as shown in Fig. 5, has alleviated the issue of sudden divergence.